Circular No. 6026

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Further reports, received in part via the SL9 message center, indicate

considerable confusion concerning the detection of the impact of fragment F = 16. This had been predicted to occur 95 percent of a jovian rotation after the E impact, but because the F prediction was early this impact point almost coincided with the E impact point, which was still very active. Nevertheless, it appears that the plume from the F impact was observed at the European Southern Observatory, but only at wavelengths of 5–12 μm. B. Mosser, T. Livengood and H. U. Käufl made the detection on July 18.060 UT, and the plume remained at the limb for more than 20 min.

As expected, the impact of fragment G = 15 was the most spectacular so far, and the first report received was that from H. Nguyen at the South Pole. The University of Chicago’s South Pole Infrared Explorer (SPIREX), a 0.60-m telescope operating mainly at 2.36 μm, reported this event on July 18.320 UT. CASPIR 2.34-μm images and IRIS K-grism cubes (cf. IAUC 6024) showed an initial flash on July 18.315 (attributed to fragment G_{2} = 15b), half as bright as the C-impact site, brightening by about a factor of four on July 18.316 and then stable until 18.319, at which time the G (or G_{1} = 15a) impact saturated the detectors and produced brilliant diffraction spikes. After July 18.340 the G impact site had decreased in brightness in K to about four times that of the C impact site, but it was detectable in IRIS bands from J to M, as well as in the visible. S. Nakano, Sumoto, Japan, communicates that many amateur astronomers, some using telescopes of aperture only 0.10 m, have visually detected the G impact site; according to A. Nakamura and K. Ito the region—larger than the Great Red Spot—was very dark, with a wave to the west like that of site A.

Although some observers have suggested that there is substantial drifting of the impact sites in longitude, M. Brown, P. Nicholson and others find that the rotational velocity of the impact sites is in fact very close to that of System III (period 0.413538 day). As further impacts occur, and as many of the impact sites persist, confusion is likely to mount. The following tabulation gives the first central-meridian passage on July 18 UT for each impact site as derived from Nakano’s predictions and adjusted (with a constant shift) to recent observations of sites A, C, E and G: site G, July 18.017 UT; D, 18.025; S, 18.028; R, 18.035; Q, 18.056; B, 18.070; N, 18.070; H, 18.101; F, 18.138; T, 18.145; E, 18.161; V, 18.161; A, 18.199; C, 18.241; P, 18.269; K, 18.307; U, 18.311; W, 18.314; L, 18.386.

1994 July 18 Brian G. Marsden